OpenStudy (katherinesmith):

how many degrees are in 3(pi)/2 radians?

OpenStudy (anonymous):

are you just trying to convert?

OpenStudy (katherinesmith):

\[\frac{ 3 }{ 2}\]

OpenStudy (katherinesmith):

my pi symbol didn't show up but i guess so

OpenStudy (anonymous):

from radians to degrees?

OpenStudy (katherinesmith):

\[\frac{ 3\pi }{ 2 }\]

OpenStudy (psymon):

Oops, 180/pi

OpenStudy (katherinesmith):

i guess so. the question is the only information i got.

OpenStudy (anonymous):

you divide 180 by 2

OpenStudy (katherinesmith):

so 90

OpenStudy (anonymous):

then multiply by

OpenStudy (psymon):

\[\frac{ 3\pi }{ 2 }*\frac{ 180 }{ \pi } \]

OpenStudy (anonymous):

by 3

OpenStudy (anonymous):

so 180/ 2 then multiply that by 3 and that's your answer in degrees

OpenStudy (anonymous):

so answer is 270 degrees

OpenStudy (psymon):

radians -> degrees, always is multiply what you have by 180/pi degrees - > radians, always multiply by pi/180.

OpenStudy (anonymous):

you haeve to divide first though

OpenStudy (psymon):

Its just built into the formula. You multiply and divide either way. Could easily multiply first if ya had to. Either way, the formula is the formula, just gotta use it however you like, lol.

OpenStudy (anonymous):